OpenStudy (anonymous):

I just started calculus and I need help solving lim x->0 sin x/ 2x^2-x

OpenStudy (cwrw238):

if we substitute x = 0 into this function we get 0 / 0 which is indeterminate use l'hopitals rule find derivative of top and bottom = cos x / 4x - 1 limit = 1 / -1 = -1

OpenStudy (noelgreco):

Factor out a sin(x)/x That's one of the first limits taught. Then use the rules of limits to solve.

OpenStudy (cwrw238):

ah - yes - thats a better way to do it

OpenStudy (anonymous):

how would you factor out a a limit?

OpenStudy (noelgreco):

Factor the denominator to x(2x-1)

OpenStudy (cwrw238):

you won't you have done l'hopitals rule yet i don't expect

OpenStudy (anonymous):

i got that part. i just don't know what the rest is

OpenStudy (noelgreco):

\[\lim_{x \rightarrow 0}\frac{ \sin x }{ x } \times \lim_{x \rightarrow 0}\frac{ 1 }{ 2x-1 }\]

OpenStudy (anonymous):

what comes after that? or is that it?

OpenStudy (noelgreco):

Do you know what the first limit is? - like I said, that's quite important and is always taught early-on in beginning calculus.

OpenStudy (noelgreco):

http://www.youtube.com/watch?v=Ve99biD1KtA

OpenStudy (noelgreco):

BTW, Cindy, you don't need to know that whole proof, but MEMORIZE \[\lim_{x \rightarrow 0}\frac{ \sin x }{ x }=0\] You'll be happy you spent the gray cells.

OpenStudy (anonymous):

Thank you for that!